OpenStudy (anonymous):

Simplify (6 +i)(8 – 3i) 51 – 10i 48 – 32i 45 – 10i 48 – 10i – 3i^2

OpenStudy (angela210793):

48-18i+8i+3=51-10i

OpenStudy (anonymous):

Assuming i is the square root of -1 \[(6+i)(8-3i) = 48-18i+8i-3i^2\] \[= 48-10i+3\] \[= 51-10i\]